find three consecutive integers such that four times their sum is equal to the product of the larger two numbers.
Answer by mananth(16949)

let the integers be x,x+1,x+2
4(x+x+1+x+2)=(x+1)(x+2)
4(3x+3)=(x+1)(x+2)
12x+12=x^2+3x+2
rearrange
x^2-9x-10=0
x^2-10x+x-10=0
x(x-10)+1(x-10)=0
(x-10)(x+1)=0
x=10 OR x=-1
10,11,12
-1,0,1